How many miles of CO gas are contained in a 1.5L container at 1.90atm and 45°C
1
Expert's answer
2020-03-18T12:55:58-0400

According to ideal gas state equation:

"p*V=n*R*T"

p is pressure in Pa = 1.90 atm*101325 Pa/atm = 192517.5 Pa

V is volume in m^3 = 1.5 l*0.001 m^3/l = 0.0015 m^3

n is moles

R is gas constant = 8.31 J/(K*mole)

T is temperature in K = 45 + 273.15 = 318.15 K

Hence,

"n=p*V\/R\/T=192517.5*0.0015\/8.31\/318.15=0.109"

Answer: 0.109 moles